OpenStudy (anonymous):

The population of a city (in millions of people) is modled by p(t_ = 3e^t/5 with t=0 representing the year 2000. a. When(to the nearest year) will the population be 100 million? when will it be a billion b. At what rate is the population increasing in 2005? IN 2010? Indicate units. c. Use the 2nd derivative to justify why p'(10)>p'(5)

OpenStudy (campbell_st):

quick question is the model 1. \[P(t) = 3e^{\frac{t}{5}}\] or 2. \[P(t) = \frac{3e^t}{5}\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

a. P(t) is equal to 100 in the first part. 100 = 3e^(t/5) We're solving for t. 100/3 = e^(t/5) Take natural log of both sides, which will get rid of the e log (100/3) = t/5 Multiply both sides by 5 t = 5log(100/3) Use calculator, and its 7.614, or 2000 + 8, to the nearest years: 2008 Try the second part yourself.

OpenStudy (anonymous):

okay but how do you do part c.